OpenStudy (anonymous):

(sqrt(2 x+3)=2+sqrt(x-2), {x})

OpenStudy (raden):

3, just guess...

OpenStudy (raden):

first, if u squaring to both sides. what u get ?

OpenStudy (raden):

do u understand this @getexcool ?

OpenStudy (anonymous):

yeah i understand. I am working on it now

OpenStudy (raden):

ok, tell me what did u get now

OpenStudy (raden):

(sqrt(2x+3))^2 = ... ? (2 + sqrt(x-2))^2 = .... ?

OpenStudy (raden):

hints : (sqrt(a))^2 = a (a+b)^2 = a^2 + 2ab + b^2

OpenStudy (raden):

btw, my guess is still incomplete.. i hope u can find the answer completely :)

OpenStudy (anonymous):

thanks